Two point charges, +3.63 µC and -6.15 µC, are separated by 1.26 m. What is the electric potential midway between them?
Answer #1

U1 = kq1 / (d/2) = 2k q1 / d <<<<<<<<

U2 = kq2 / (d/2) = 2k q2 / d <<<<<<<<

so U total = U1 + U2 = (2k/d) (q1 + q2) , d = 1.26 m

=2*9.0x10^9* ( 3.63x10^-6C - 6.15x10^-6)/1.26m) =

you should include the sign when finding the potential .

potential becomes -3.6*10^4